### 26 Feb 2024 v3

# Ricci flow-guided autoencoders in learning time-dependent dynamics

Abstract
We present a manifold-based autoencoder method for learning nonlinear dynamics in time, notably partial differential equations (PDEs), in which the manifold latent space evolves according to Ricci flow. This can be accomplished by simulating Ricci flow in a physics-informed setting, and manifold quantities can be matched so that Ricci flow is empirically achieved. With our methodology, the manifold is learned as part of the training procedure, so ideal geometries may be discerned, while the evolution simultaneously induces a more accommodating latent representation over static methods. We present our method on a range of numerical experiments consisting of PDEs that encompass desirable characteristics such as periodicity and randomness, remarking error on in-distribution and extrapolation scenarios. Our code is available at https://github.com/agracyk2/Ricci-flow-guided-autoencoders-in-learning-time-dependent-dynamics.
1 Introduction
Data-driven techniques, fueled by machine learning, have become a useful tool in exploring partial differential equations (PDEs), such as in instances of finding solutions. A variety of frameworks have been developed with the purpose of finding approximations to time-dependent PDE solutions including encoder-decoder methods. Encoder-decoder methods are distinguished by a series of neural networks in which there exists an intermediate encoded space stage, which can be variously characterized, that is decoded to produce a PDE solution output. Geometric dynamic variational autoencoders (GD-VAEs) have shown great promise in learning disrcetized time-dependent dynamics (Lopez and Atzberger, 2022), which are highlighted by an underlying latent space with either geometric features or dynamics in the latent space itself. We focus on an encoder-decoder method, emphasizing a low-dimensional manifold latent space which is imposed with dynamical evolution, that also holds the advantages that GD-VAEs achieve, including competitive error, generalization, extrapolation ability, and robustness.
Examples of existing manifold-based encoder methods involve scenarios of data mapped to Euclidean space without structure, in which a projection onto a manifold is subsequently taken for decoding. This in itself requires an optimization problem to be solved per every iterate of the training procedure. Additionally, fixed geometries are possibly less than ideal for their compatibility with the PDE data, observable from an error perspective. Furthermore, potential trial and error is needed to determine optimal latent structure, if found at all. Furthermore, evolution in the latent space has shown competitive performance over static methods (Lopez and Atzberger, 2022). We propose a method in which no optimization per iterate is necessary, which incorporates manifold geometries with such evolution.
A geometric flow, the foundation of our method, is a tool to perform dynamics among the manifold, in which we choose Ricci flow. We first represent the initial PDE data in a domain to parameterize the manifold, which is then encoded directly to a point on the manifold. The points that compose this manifold evolve according to such a flow, in which we use a physics-informed neural network to simulate the evolution, matching quantities among the manifold to those that solve the physics-informed flow. Using neural networks to learn the manifold directly, we can learn implicit geometric features that accommodate the PDE data to an unmatched degree over geometries that are predetermined.
Ricci flow is an example of an intrinsic geometric flow, meaning the relevant quantities computed in the PDE flow are inherent to the manifold itself, and so the manner in which it is embedded in space is not involved in the PDE computation and is otherwise irrelevant. An extrinsic geometric flow is a flow equation in which the embedding impacts the informational quantities used in the PDE, and so the embedding does indeed matter and is factored into how the manifold evolves subject to the flow. The metric, a quantity inherent to the manifold, can be computed as an inner product of the tangent vectors (Millman and Parker, 1977). The location of these tangent vectors can be arbitrarily changed in space in which the inner product is unchanged, assuming the relative positioning and scaling of these vectors is unaffected, making a PDE consisting of metric quantities an intrinsic flow.
Our method, captured by an intrinsic flow, is highlighted by this fact. This allows the manifold to displace through space, aside from the movements due to the flow, during its period of evolution. The embedding is irrelevant, and can change through the training procedure to maximize effectiveness in solving the objective function.
The intermediary stage of the parameterization domain acts similarly to the unstructured data in Euclidean space in the GD-VAE setting. These outputs are given heightened flexibility and permitted to lie in unrestricted subdomains before subsequently being mapped to the manifold. In this sense, our method applies no greater restrictions upon the latent space by using this middle step of unstructured data of the manifold’s parameterization, the only of which being this domain is one dimension lower than that of the embedding in which the unstructured data in the GD-VAE case is contained.
While the implicitly-learned latent space captured by an intrinsic flow allows high precision in the task of replicating data from the same underlying distribution as seen as part of the training procedure, a challenge resides in achieving precision in areas such as generalization of the data and extrapolation of data with fundamental qualities of those seen in training, but sufficiently different to be categorized as data that is extended outside of the training procedure. The manifold latent space offers promise in extending the autoencoder baseline to one that yields greater ability to infer dynamics from never-before-seen data, as well as incorporate features of the data into such a structure (Lopez and Atzberger, 2022).
2 Problem formulation
Problem setting. The task of interest is to learn the solution of a parameterized partial differential equation of the form
$\partial_{t}\phi+\mathcal{D}_{\alpha}[\phi]=0,$ (1)
subject to initial data $\phi(\cdot,0)=\phi_{0}$, where $\phi:\mathcal{X}\times[0,T]\rightarrow\mathbb{R}$ is the solution, and $\mathcal{D}_{\alpha}$ is a differential operator with parameter $\alpha$. $\alpha$ may depend on evaluations over the input space $\mathcal{X}$ as well as time $t\in[0,T]$. In particular, we seek a mapping
$(\phi_{0},t)\in C(\mathcal{X};\mathbb{R})\times[0,T]\rightarrow\phi_{t}\in C(\mathcal{X};\mathbb{R}).$ (2)
We devise an autoencoder framework, characterized by a latent space that consists of collections of points that lie directly along a continuum of manifolds $\mathcal{M}_{\hat{\tau}}$ embedded in Euclidean space $\mathbb{R}^{d}$, $\hat{\tau}\in[0,\tau]$ for some $\tau$, to learn $\phi_{t}$.
Ricci flow. Ricci flow is a natural framework for accomplishing goals of incorporating input data features into manifold structure and generalization capability while allowing high precision in the task of producing the output. Ricci flow unites the separate features of an evolving latent space and a manifold latent space into a single scheme. Ricci flow is a partial diffferential equation that describes changes in the Riemannnian metric. Let $g_{ij}:\mathcal{U}^{1}\times\mathcal{U}^{2}\times\ldots\times\mathcal{U}^{m}\times[0,\tau]\rightarrow\mathbb{R}$ be the components of the Riemmannian metric matrix $g$ over the manifold parameterization domain $\mathcal{U}^{1}\times\mathcal{U}^{2}\times\ldots\times\mathcal{U}^{m}=\mathcal{U}$ and a time domain $[0,\tau]$. Ricci flow is the metric evolution equation
$\partial_{t}g(u,t)=-2\text{Ric}(g(u,t)),$ (3)
for $u\in\mathcal{U},t\in[0,\tau]$, such that the metric solved by 3 reflects that of the manifold $\mathcal{M}_{\hat{\tau}}$ over its respective time interval. The solution to this PDE can be regarded as an $m\times m$ matrix, $g(u,t)$, with components $g_{ij}$. The right-hand term $\text{Ric}(g(u,t))$ is the Ricci tensor. Ricci flow is similar to a geometric diffusion equation in which curvature is uniformized, and geometries contract or expand.
Metric coefficients. The coefficients of the Riemannian metric, $g_{ij}$, can be computed using the tangent vectors to the manifold at a location corresponding to a point in the manifold parameterization domain and time domain $\mathcal{U}\times[0,\tau]$. The coefficients of the metric can be computed at general times as (Millman and Parker, 1977)
$g_{ij}(u,t)=\langle\partial_{i}\mathcal{E}(u,t),\partial_{j}\mathcal{E}(u,t)\rangle,$ (4)
where $\partial_{\ell}$ is the component-wise partial derivative with respect to the $\ell$-index in $u=(u^{1},u^{2},\ldots,u^{m})$ belonging to the parameterization domain $\mathcal{U}$, and $\mathcal{E}:\mathcal{U}\times[0,\tau]\rightarrow\mathbb{R}^{d}$ is the function mapping to the manifold continuum. The Ricci flow PDE is primarily governed based on initial coefficients $g_{ij,0}$, but in our methodology, it of interest to compute the metric coefficients among arbitrary times $\tilde{\tau}_{i}\in U([0,\tau])$ in a uniform collocation procedure. The behavior of the manifold is governed by the change of the metric coefficients under Ricci flow such that equation 4 holds.
Ricci tensor. Given explicit knowledge of the Riemannian metric coefficients, one may compute the Ricci tensor using Christoffel symbols and the Riemannian curvature tensor. The Christoffel symbols are a collection $\{{\Gamma_{ij}}^{l}\}_{i,j,l},1\leq i,j,l\leq d-1$ evaluated at particular $(u_{i},t_{i})$, given by (Millman and Parker, 1977)
${\Gamma_{ij}}^{l}=\frac{1}{2}\sum_{k=1}^{d-1}g^{kl}(\partial_{j}g_{ik}-\partial_{k}g_{ij}+\partial_{i}g_{kj}).$ (5)
$g^{kl}$ denote the elements of the inverse matrix of $g$ at such a $(u_{i},t_{i})$. The Riemannian tensor is a collection $\{R_{ijk}^{l}\}_{i,l,j,k},1\leq i,l,j,k\leq d-1$, which can be computed as (Millman and Parker, 1977)
$R_{ijk}^{l}=\frac{\partial}{\partial u^{j}}{\Gamma_{ik}}^{l}-\frac{\partial}{\partial u^{k}}{\Gamma_{ij}}^{l}+\sum_{p=1}^{d-1}({{\Gamma}_{ik}}^{p}{\Gamma_{pj}}^{l}-{\Gamma_{ij}}^{p}{\Gamma_{pk}}^{l}).$ (6)
Finally, the Ricci tensor is a contraction of the Riemannian tensor, and so
$\text{Ric}(g(u,t))_{ik}=R_{ilk}^{l}=\sum_{l=1}^{d-1}R_{ilk}^{l},$ (7)
where the term in the middle uses Einstein notation, is the corresponding Ricci tensor at a particular $(u,t)$, which is equivalently an analog of the trace of the Riemannian tensor.
3 Methodology
We present Ricci flow-guided autoencoders to learn PDE dynamics. From a collection of PDE observations with known initial data with parameter $\alpha$, we develop a series of neural networks, first mapping to a parameterization domain, then a manifold encoding, which is then decoded as a prediction to PDE output. An interposed neural network acts as the metric as well as a solution to Ricci flow, in which the relation between the tangent vectors to the manifold and the metric acts as Ricci flow enforcement.
Figure 1: We illustrate our methodology. Initial PDE data is mapped to a parameterization domain ($u$), then encoded onto the manifold ($\mathcal{E}_{\theta_{\mathcal{E}}}$), then decoded ($\mathcal{D}_{\theta_{\mathcal{D}}}$) as the PDE solution. The metric neural network $g_{\theta_{g}}$ is subject to Ricci flow, with an enforcing constraint.
Neural network design. Necessary PDE information acts as input training data into a parameterization neural network $\mathcal{P}_{\theta_{\mathcal{P}}}$, mapping such data to a point in the parameterization domain $\mathcal{U}$ of the manifold continuum $\mathcal{M}_{\hat{\tau}}$. In particular, we construct a neural network $\mathcal{P}:\widetilde{\Phi}\times\Theta_{\mathcal{P}}\rightarrow\mathcal{U}$, where $\widetilde{\Phi}$ is an $N$-dimensional vector containing data of the PDE initial conditions $\{\phi_{0}^{i}\}_{i}$, $\phi_{0}^{i}=\{(\phi_{0}^{i}(x_{1},0),\ldots,\phi_{0}^{i}(x_{N},0)):x_{j}\in\Omega\}$ for mesh $\Omega$. We denote $\mathcal{P}_{\theta_{\mathcal{P}}}(\phi)=\mathcal{P}(\phi,\theta_{\mathcal{P}})$. The purpose of this neural network, instead of directly mapping PDE information into a manifold latent space, is so that the partial derivatives $\partial_{i}\mathcal{E}(u,t)$, hence tangent vectors, of the manifold may be taken with respect to the same domain in which the metric $g$ is defined. Without this, 4 cannot be formulated, which invalidates one of the terms we will develop in our loss function.
We propose a physics-informed neural network (PINN) framework for finding the metric solution $g$ to the Ricci flow equation. The PINN $g_{\theta_{g}}$ is constructed via standard neural network architecture, but is substituted into the pertinent PDE and is subsequently minimized as a PDE residual, as this is set to 00 in what should be satisfied (Raissi et al., 2017). Differentiation of the PINN can be performed using automatic differentiation techniques found in deep learning libraries. To construct the metric, let $g:\mathcal{U}\times[0,\tau]\times\Theta_{g}\rightarrow\mathbb{M}^{m\times m}$, where $\mathbb{M}^{m\times m}$ is the set of $m\times m$ matrices. Generally, $m=d-1$. Since the metric should satisfy Ricci flow, the PDE residual to be minimized in our PINN setup is $\partial_{t}g_{\theta_{g}}+2\text{Ric}(g_{\theta_{g}})$ (Jain et al., 2022).
Lastly, we construct encoder and decoder networks. The encoder maps a point along the learned parameterization domain directly onto the manifold embedded in Euclidean space at a specified time, and so the encoder additionally has a temporal input parameter corresponding to a state of evolution of the manifold. The encoder is given by $\mathcal{E}:\mathcal{U}\times[0,\tau]\times\Theta_{\mathcal{E}}\rightarrow\mathcal{M}_{\hat{\tau}}$, for $\hat{\tau}\in[0,\tau]$. The decoder, $\mathcal{D}:\mathcal{M}_{\hat{\tau}}\times\Theta_{\mathcal{D}}\rightarrow\mathbb{R}$, maps the point in Euclidean coordinates directly to known PDE solutions. A series of compositions of functions is taken to yield the final autoencoder output, with loss restrictions imposed on the various functions at the intermediate steps to ensure the desired conditions, such as the manifold evolving under Ricci flow, are satisfied.
Loss function. We construct our training objective function as follows. {strip}
$\theta$ $=\operatorname*{arg\,min}_{\theta\in\Theta}\mathcal{L}_{Ric}(\theta)+\lambda_{dec}\mathcal{L}_{dec}(\theta)+\lambda_{met}\mathcal{L}_{met}(\theta),$ (8)
$\mathcal{L}_{Ric}(\theta)$ $=\operatorname{\mathbb{E}}_{t}\operatorname{\mathbb{E}}_{\Phi}\Big{[}\frac{1}{|\mathcal{U}|^{2}}||\partial_{t}g_{\theta_{g}}(u,\tilde{\tau})+2\text{Ric}(g_{\theta_{g}}(u,\tilde{\tau}))||_{F}^{2}\Big{]},$ (9)
$\mathcal{L}_{dec}(\theta)$ $=\operatorname{\mathbb{E}}_{t}\operatorname{\mathbb{E}}_{\Phi}\Big{[}\frac{1}{N}||\mathcal{D}_{\theta_{\mathcal{D}}}(\mathcal{E}_{\theta_{\mathcal{E}}}(u,\hat{\tau}))-\phi_{t}||_{2}^{2}\Big{]},$ (10)
$\mathcal{L}_{met}(\theta)$ $=\operatorname{\mathbb{E}}_{t}\operatorname{\mathbb{E}}_{\Phi}\Big{[}\frac{1}{|\mathcal{U}|^{2}}||g_{\theta_{g}}(u,\tilde{\tau})-(J\mathcal{E}_{\theta_{\mathcal{E}}}(u,\hat{\tau}))^{T}(J\mathcal{E}_{\theta_{\mathcal{E}}}(u,\hat{\tau}))||_{F}^{2}\Big{]}.$ (11)
We use notation $J\mathcal{E}_{\theta_{\mathcal{E}}}$ to denote the Jacobian of vector $\mathcal{E}_{\theta_{\mathcal{E}}}$ with respect to $u$, i.e. $(J\mathcal{E}_{\theta_{\mathcal{E}}}(u,\hat{\tau}))^{T}(J\mathcal{E}_{\theta_{\mathcal{E}}}(u,\hat{\tau}))$ is the matrix of inner products
$(J\mathcal{E}_{\theta_{\mathcal{E}}}(u,\hat{\tau}))^{T}$ $(J\mathcal{E}_{\theta_{\mathcal{E}}}(u,\hat{\tau}))=\begin{pmatrix}\psi_{11}&\psi_{12}&\dots\\ \psi_{21}&\psi_{22}&\dots\\ \vdots&\vdots&\ddots\end{pmatrix},$ (15)
$\psi_{ij}$ $=\langle\partial_{i}\mathcal{E}_{\theta_{\mathcal{E}}}(u,\hat{\tau}),\partial_{j}\mathcal{E}_{\theta_{\mathcal{E}}}(u,\hat{\tau})\rangle,$ (16)
$\psi_{ii}$ $=||\partial_{i}\mathcal{E}_{\theta_{\mathcal{E}}}(u,\hat{\tau})||_{2}^{2}.$ (17)
We motivate our loss so that the relation between the metric and the inner product of derivatives of 4 is satisfied under Ricci flow. A difference between decoder prediction and PDE solution is minimized to ensure a proper output.
The first expectation runs over time, and the second runs over training data, taking data $\phi_{0}^{i}\sim\Phi$ for PDE data belonging to distribution $\Phi$. We use notation $|\mathcal{U}|=\dim(\mathcal{U})$. Denote $u=\mathcal{P}_{\theta_{\mathcal{P}}}(\phi_{0})$ a value within $\mathcal{U}$, which is learned via neural network $\mathcal{P}_{\theta_{\mathcal{P}}}$ and not set beforehand, corresponding to points in the training data. To clarify, $\phi_{t}=\phi(\cdot,t)$. Let $\Theta=(\Theta_{\mathcal{P}},\Theta_{g},\Theta_{\mathcal{E}},\Theta_{\mathcal{D}})$. Let $\hat{\tau}\in U([0,\tau]),\tilde{\tau}\in U([0,\tau^{\prime}])$ be uniformly sampled via a collocation procedure, where $\tau=CT$ for some constant $C$, and $\hat{\tau}$ and $\tilde{\tau}$ are times associated with each other by some linear scaling. We will generally choose $\hat{\tau}=\tilde{\tau}$. We scale the times belonging to Ricci flow $\hat{\tau}$, primarily to ensure no singularities develop, which are possible under an identity relationship. $||\cdot||_{F}$ is the Frobenius norm. All derivatives are computed via automatic differentiation found in standard deep learning libraries. The analytic expectations can be taken using empirical averages during the training procedure, additionally divided into batches with parameters updated with a stochastic optimizer.
Algorithm 1 Training algorithm
Input: Initial conditions $\{\phi_{0}^{i}\}_{i}$, solutions $\{\phi_{t_{j}}^{i}\}_{i,j}$ evaluated at $(\cdot,t_{j})\in\Omega\times[0,T]$
Setup: Set $\hat{\tau}_{i}=C_{t}t_{i}$, sample $\tilde{\tau}_{i}\in U([0,\tau])$ for some final $\tau$. Select batch size $n$
1:while $\mathcal{L}(\theta)$ is not converged do
2:     Sample $n$ initial conditions $\{\phi_{0}^{i}\}_{i\in\mathcal{I}_{n}}$ and $n$ solutions $\{\phi_{t_{j}}^{i}\}_{i\in\mathcal{I}_{n},j\in\mathcal{J}_{n}}$
3:     $u_{i}=\mathcal{P}_{\theta_{\mathcal{P}}}(\phi_{0}^{i})\in\mathcal{U}$
4:     Compute encoded point on manifold $\mathcal{E}_{\theta_{\mathcal{E}}}(u_{i},\hat{\tau}_{i})$
5:     Compute decoded solution $\mathcal{D}_{\theta_{\mathcal{D}}}(\mathcal{E}_{\theta_{\mathcal{E}}}(u_{i},\hat{\tau}_{i}))$
6:
7:Ricci flow loss
8:  Compute:
9:     $\partial_{t}g_{\theta_{g}}(u_{i},\tilde{\tau}_{i})$triangleright time derivative
10:     $g^{kl}\in G^{-1}$triangleright inverse matrix of metric
11:     $\partial_{i}g_{jk}$triangleright partial derivatives
12:     $\partial_{il}g_{jk}$triangleright second order partial derivatives
13:     ${\Gamma_{ij}}^{l}=\frac{1}{2}\sum_{k}g^{kl}(\partial_{j}g_{ik}-\partial_{k}g_{ij}+\partial_{j}g_{kj})$triangleright Christoffel symbols
14:     $R_{ijk}^{l}=\partial_{j}{\Gamma_{ik}}^{l}-\partial_{k}{\Gamma_{ij}}^{l}+\sum_{p}({{\Gamma}_{ik}}^{p}{\Gamma_{pj}}^{l}-{\Gamma_{ij}}^{p}{\Gamma_{pk}}^{l})$triangleright Riemannian tensor
15:     $\text{Ric}(g(u,t))_{ik}=\sum_{l}R_{ilk}^{l}$triangleright Ricci tensor
16:     Loss $\mathcal{L}_{Ric}(\theta)$
17:
18:Metric loss
19:
20:     Compute $\langle\partial_{j}\mathcal{E}_{\theta_{\mathcal{E}}}(u_{i},\hat{\tau}_{i}),\partial_{k}\mathcal{E}_{\theta_{\mathcal{E}}}(u_{i},\hat{\tau}_{i})\rangle$triangleright inner product constraint
21:     Formulate $\mathcal{L}_{met}(\theta)$ from above and physics-informed metric $g_{\theta_{g}}$
22:
23:Decoder loss
24:
25:     Compute $\mathcal{L}_{dec}(\theta)$ directly from decoder and $\{\phi^{i}(\cdot,t_{j})\}$
26:
27:Minimize objective.
28:
29:     Compute $\mathcal{L}_{Ric}(\theta)+\lambda_{dec}\mathcal{L}_{dec}(\theta)+\lambda_{met}\mathcal{L}_{met}(\theta)$
30:     Update $\theta$
31:end while
The loss terms are organized as a Ricci flow physics-informed term $\mathcal{L}_{Ric}$, quadratic loss $\mathcal{L}_{dec}$ to ensure the decoder matches the PDE solution, and a constraint term $\mathcal{L}_{met}$ to match the manifold evolution with the metric satisfying Ricci flow. $\lambda_{dec}$ and $\lambda_{met}$ are scaling coefficients as needed. In the final term $\mathcal{L}_{met}$, $g_{\theta_{g}}$ is the matrix of the physics-informed metric term $g_{\theta_{g}}$, where the components are summed with respect to each inner product $\langle\partial_{j}\mathcal{E}_{\theta_{\mathcal{E}}}(u_{i}),\partial_{k}\mathcal{E}_{\theta_{\mathcal{E}}}\rangle$ to enforce loss for each entry.
We will generally restrict our attention to the case $\mathcal{U}\subseteq\mathbb{R}^{2}$. as this exhibits computational simplicity in the training procedure, yet sufficient latent representation. $\mathcal{U}$ generally parameterizes a $2$-dimensional manifold, i.e. $\mathcal{E}(u,t)$ is embedded in $\mathbb{R}^{3}$.
Special cases. Special cases of Ricci flow may be under consideration, which are favorable for reasons such as lower computational cost or higher-dimensional encodings. In these scenarios, information about the manifold is set beforehand, such as the metric. Ideal latent geometries are sacrificed at the gain of some other benefit. We discuss special cases in Appendix D. Coordinate transformations may be appropriate under select special cases. Our method in such a setting is discussed in Appendix C.
4 Experiments
We present Ricci flow-guided autoencoders on a series of numerical experiments for approximate solutions of PDEs given collections of known data. We compare our method to existing GD-variational autoencoder methods in (Lopez and Atzberger, 2022), as well as other forms of autoencoder baselines. Results for both the Ricci flow method and alternative methods are collected using identical initial condition input spaces. The GD-VAE method is performed using an evolution mapping $\phi_{t}\rightarrow\phi_{t+\tilde{t}}$, where $\tilde{t}$ is some time-increment used to progress the solution to a later time. Predictions are retrained for new values of $\tilde{t}$.
The highlighting feature of our method is the manifold latent space subject to Ricci flow. We will reserve comparisons for recreating data as seen as part of the training procedure for manifold-based methodology to emphasize results for such a variety of framework.
[…]
light-gray 1-d Burger’s equation
method
$t=0$
$t=0.25$
$t=0.5$
$t=0.75$
$t=1$
[…]
LightCeruleanRicci flow
$1.76\pm 1.38$
$1.53\pm 0.502$
$1.43\pm 0.586$
$1.38\pm 0.740$
$2.19\pm 0.970$
GD-VAE
$14.4\pm 18.9$
$4.52\pm 5.47$
$4.90\pm 6.88$
$3.75\pm 3.68$
AE, g-proj.
$10.3\pm 10.6$
$10.4\pm 7.89$
$5.29\pm 11.9$
$2.24\pm 1.13$
[…]
light-gray 1-d diffusion-reaction
$t=0$
$t=0.25$
$t=0.5$
$t=0.75$
$t=0.9$
[…]
LightCerulean Ricci flow
$0.274\pm 0.136$
$0.184\pm 0.0755$
$0.215\pm 0.0932$
$0.429\pm 0.199$
$1.14\pm 0.410$
GD-VAE
$2.10\pm 2.81$
$2.94\pm 6.17$
$1.53\pm 0.886$
$2.24\pm 0.881$
AE, g-proj.
$1.78\pm 2.59$
$1.28\pm 0.903$
$1.41\pm 0.595$
$2.77\pm 1.33$
[…]
light-gray 1-d diffusion-reaction extrapolation
$\mathcal{A}_{2}^{new,1},t=0.35$
$\mathcal{A}_{2}^{new,2},t=0.35$
$\mathcal{A}_{2}^{new,3},t=0.35$
[…]
LightCerulean Ricci flow
$10.4\pm 15.4$
$17.8\pm 9.62$
$30.8\pm 13.1$
GD-VAE
$14.9\pm 17.0$
$21.9\pm 13.9$
$36.1\pm 18.6$
AE
$19.4\pm 27.0$
$23.6\pm 12.8$
$47.5\pm 26.3$
[…]
light-gray 2-d Navier-stokes equation
$t=0$
$t=0.25$
$t=0.5$
$t=0.75$
$t=1.0$
[…]
LightCerulean Ricci flow
$4.44\pm 1.48$
$3.97\pm 1.57$
$5.26\pm 1.18$
$8.46\pm 1.09$
$9.25\pm 1.53$
GD-VAE
$84.3\pm 8.92$
$88.4\pm 12.8$
$93.5\pm 5.05$
$92.1\pm 4.02$
Table 1: We compare our methodology to those existing, reporting relative $L^{1}$ errors on 30 testing sets. All errors are taken $10^{-2}$. Times corresponding to the increment $\tilde{t}=0$ are omitted for comparisons since this is a trivial learning task in these frameworks. Such error by a decrease in order of magnitude for the diffusion-reaction experiment has partial attribution to the enhanced MLP architecture (Appendix F).